Three times the age of francis is four more than the age of sally. In three years the sum of their ages will be 30. Find their present ages
Answer by stanbon(75887)

Three times the age of francis is four more than the age of sally. In three years the sum of their ages will be 30. Find their present ages
------
3F = S + 4
F+3 + S+3 = 30
--------------------
S = 3F-4
S + F = 24
--------------------
3F-4 = 24-F
4F = 28
F = 7 (Francis's age)
S = 3F - 4 = 17 (Sally's age)
--------------
Cheers,
Stan H.
-------------